OpenStudy (anonymous):

helllllllllllp plsssssss The complex number z is given by z = (√3) + i. Showing your working, express in the form x + iy, where x and y are real, (a) iz*/z

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@nincompoop

OpenStudy (irishboy123):

do you know what z* means or is?

OpenStudy (anonymous):

conjugate

OpenStudy (anonymous):

@phi

OpenStudy (phi):

write down what you have so far

OpenStudy (anonymous):

first tell me should we multiply like this i(\[\sqrt{3}-i\] )= sqrt3i -i^2

OpenStudy (phi):

yes, that is correct. however, remember i*i is -1

OpenStudy (anonymous):

yes but am stuck.. am not getting the right answer -_-

OpenStudy (anonymous):

could solve pls

OpenStudy (phi):

your not stuck. just not finished. so fare you have \[ \frac{i \ z*}{z} = \frac{1+ \sqrt{3} \ i}{\sqrt{3} + i} \]

OpenStudy (phi):

to get rid of a complex number in the denominator, multiply it by its complex conjugate. to keep things equal, do the same to the top

OpenStudy (anonymous):

i already done all this... am stuck.. i think i have not done it right

OpenStudy (anonymous):

let me try again

OpenStudy (phi):

first, what do you get for \[ (\sqrt{3} + i)(\sqrt{3} - i)\]?

OpenStudy (anonymous):

its the signs problem

OpenStudy (phi):

do you know FOIL? for multiplying

OpenStudy (anonymous):

no

OpenStudy (phi):

FOIL is a way to remember what to multiply \[ (\sqrt{3} + i)(\sqrt{3} - i) \] F irst \( \sqrt{3} \sqrt{3} \) Outer \( \sqrt{3} \cdot -i \) Inner \( i \cdot \sqrt{3} \) Last \( i \cdot -i\)

OpenStudy (phi):

the other way to remember is use the distributive law

OpenStudy (anonymous):

am having a problem with the numerator

OpenStudy (phi):

First, what do you get for the bottom ?

OpenStudy (anonymous):

\[(1+\sqrt{3}i) \times (\sqrt{3}-i)\]

OpenStudy (anonymous):

for denominator ive got 4

OpenStudy (phi):

ok. now the top. You can use FOIL \[ (1+\sqrt{3}i) (\sqrt{3}-i) \] First \( 1 \cdot \sqrt{3} \) Outer \( 1 \cdot -i\) Inner \( \sqrt{3} i \cdot \sqrt{3} \) Last \( \sqrt{3} i \cdot -i\)

OpenStudy (anonymous):

then what should i get ?

OpenStudy (phi):

simplify each term. Expect 2 reals (no i) and 2 imag (with an i) the first two F and O are easy to simplify

OpenStudy (anonymous):

sorry ?

OpenStudy (anonymous):

yeah i have understood and got

OpenStudy (phi):

the F (i.e. First) term is 1 * sqr(3) which simplifies to sqr(3)

OpenStudy (anonymous):

yes

OpenStudy (phi):

now simplify the next term (the Outer)

OpenStudy (anonymous):

-i

OpenStudy (phi):

ok, now the Inner term

OpenStudy (anonymous):

3i

OpenStudy (phi):

and the Last

OpenStudy (anonymous):

sqrt 3

OpenStudy (phi):

now add them up and "combine like terms" \[ \sqrt{3} -i+3i +\sqrt{3} \]

OpenStudy (anonymous):

the answer is right now.. thank you so much.. i have another problem

OpenStudy (phi):

please make it a new post. This one is a bit long.

OpenStudy (anonymous):

yes